Solve for x analytically and give the exact solution in real or complex form.
square root of 11-x = x+1.

Answer by edjones(8007)

sqrt(11-x) = x+1
11-x=x^2+2x+1 square each side
x^2+3x-10=0 subtract the left side from the right.
(x+5)(x-2)=0
x=-5, x=2
.
Ed